use elimination to solve the system x-4y=1 and x+2y=19 for x
a.-11
b.3
c.24
d.13

Answer by 303795(602)

Equation 1 : x-4y=1
Equation 2 : x+2y=19
Multiply the second equation by 2 to get
New Equation 2 : 2x+4y=38
Add the new equation 2 to the original equation 1 to get
3x = 39
x=13
Substitute x=13 into one of the equations to find the value for y(=3).